# Oriented matroids and type $\mathbb{A}$ cluster categories

### v2

@@ 3. Constructing the oriented matroids
For now, let $T$ be a triangulation of a convex $m$-gon, where $m\geqslant 4$. We construct an oriented matroid $\mathcal{M}_{T}$ from $T$ by specifying its chirotope. Given $x,y\in[m]$ with $x\neq y$, we denote the arc with vertices $x$ and $y$ by $xy$. We refer to a subset of $[m]$ of size $k$ as a $k$-subset.
@@
If the triangulation $T$ contains arcs $xy$ with $a\preccurlyeq x\prec y\prec b$, then $T$ restricts to a triangulation of the polygon with vertices $V=[m]\setminus\{v\in[m]:x\prec v\prec y\}$, where we also have $V\supseteq\{a,b,c,d\}$. Hence, we can assume that there are no such arcs $xy$ of $T$, and likewise assume that $T$ contains no arcs $xy$ with $b\preccurlyeq x\prec y\prec c$, $c\preccurlyeq x\prec y\prec d$, or $d\preccurlyeq x\prec y\prec a$. Thus, if $x$ is a vertex with $a\preccurlyeq x\prec b$, then there must be an arc (or boundary edge) $xy$ of $T$ with either $b\preccurlyeq y\prec c$, $c\preccurlyeq y\prec d$, or $d\preccurlyeq y\prec a$. If the middle case holds, then we are done, so we can assume ($\ast$) that the first or the third case holds for all $x$ with $a\preccurlyeq x\prec b$. If we have $a\preccurlyeq x\prec x^{\prime}\prec b$, with $x^{\prime}y^{\prime}$ an arc of $T$, then we cannot have $b\preccurlyeq y\prec c$ and $d\preccurlyeq y^{\prime}\prec a$, since then $xy$ and $x^{\prime}y^{\prime}$ cross. Thus, let $\overrightarrow{x}$ be the rightmost vertex with $a\prec\overrightarrow{x}\prec b$ such that there is an arc $\overrightarrow{x}\overrightarrow{y}$ of $T$ with $d\preccurlyeq\overrightarrow{y}\prec a$. Similarly, let $\overleftarrow{x}$ be the leftmost vertex with $a\prec\overleftarrow{x}\prec b$ such that there is an arc $\overleftarrow{x}\overleftarrow{y}$ of $T$ with $b\preccurlyeq\overleftarrow{y}\prec c$. By assumption ($\ast$), there are no vertices $x$ with $\overrightarrow{x}\prec x\prec\overleftarrow{x}$. In fact, one can see that $\overrightarrow{x}=\overleftarrow{x}$, since one can otherwise consider the triangle with the side $\overrightarrow{x}\overleftarrow{x}$, which must either contain an arc $\overrightarrow{x}y$ with $b\preccurlyeq y\prec c$ or an arc $\overleftarrow{x}y$ with $d\preccurlyeq y\prec a$.
We thus let $x_{ab}=\overrightarrow{x}=\overleftarrow{x}$. One can deduce the existence of analogous vertices $x_{bc}$, $x_{cd}$, and $x_{da}$ with $b\preccurlyeq x_{bc}\prec c$, $c\preccurlyeq x_{cd}\prec d$, and $d\preccurlyeq x_{da}\prec a$. It can be seen that $T$ therefore contains a quadrilateral with vertices $x_{ab}$, $x_{bc}$, $x_{cd}$, and $x_{da}$. The diagonal $xy$ of this quadrilateral which is an arc of $T$ then gives us the arc in the statement of the lemma. ∎
We now need to show that $\mathcal{M}_{T}$ is a well-defined oriented matroid. We will use [2, Corollary 3.6.3], which says that $\mathcal{M}_{T}$ is well-defined if every six-element restriction of $\mathcal{M}_{T}$ is realisable. Hence, we want to understand restrictions of $\mathcal{M}_{T}$.
@@
Given an arc $xy$ which separates $V$, set $x^{V}:=\max_{\circlearrowleft}V_{x}$ and $y^{V}:=\max_{\circlearrowleft}V_{y}$, where $\max_{\circlearrowleft}V_{x}$ is the unique element of $V_{x}$ such that $v\preccurlyeq\max_{\circlearrowleft}V_{x}\preccurlyeq x$ for all elements $v\in V_{x}$, and $\max_{\circlearrowleft}V_{y}$ is defined similarly. Define $T_{V}:=\{x^{V}y^{V}:xy\text{ an arc of $T$ which separates }V\}$.
@@
Let $\{a\prec b\prec c\prec d\}\subseteq V$. Then $\chi_{T}(a,b,c,d)=+1$ if and only if there is an arc $xy$ of $T$ with $a\prec b\preccurlyeq x\prec c\prec d\preccurlyeq y$. This then the case if and only if $a\prec b\preccurlyeq x^{V}\prec c\prec d\preccurlyeq y^{V}$. Finally, this is the case if and only if there exists an arc $x^{V}y^{V}$ of $T_{V}$ such that this holds, which is the condition for $\chi_{T_{V}}(a,b,c,d)=+1$. ∎
Hence, using [2, Corollary 3.6.3], in order to show that $\chi_{T}$ is a well-defined chirotope, it suffices to show that $\chi_{T}$ is realisable for $T$ a triangulation of a hexagon.
@@
If $T$ is the triangulation $\{26,36,46\}$, then one can check that $\chi_{T}$ is positive on all 4-subsets of $[6]$ ordered in the usual way. Hence, $\chi_{T}$ is realised by the three-dimensional cyclic polytope with six vertices [2, Section 9.4]. By rotating, we have that $\chi_{T}$ is realisable for any triangulation $T$ of the form $\{(a-4)a,(a-3)a,(a-2)a\}$.
@@
Hence, we obtain the following result by applying [2, Corollary 3.6.3].
@@
First, note that every 5-subset of $[m]$ is a circuit of $\mathcal{M}_{T}$, since every 4-subset is a basis. Thus, let $\underline{C}=\{a\prec b\prec c\prec d\prec e\}\subseteq[m]$ be the underlying set of a signed circuit of $\mathcal{M}_{T}$, where $a\prec b\prec c\prec d\prec e$. Clearly the signs of $C$ are determined by the restriction $\mathcal{M}_{T}(\underline{C})=\mathcal{M}_{T_{\underline{C}}}$. We have that $T_{\underline{C}}$ is a triangulation of a pentagon. Up to rotation, there is only one triangulation of a pentagon, so we can rotate and assume that $T_{\underline{C}}=\{be,ce\}$. We then have that $\mathcal{M}_{T_{\underline{C}}}$ is then the oriented matroid of a three-dimensional cyclic polytope with five vertices. This implies that $C=(\{a,c,e\},\{b,d\})$ or $C=(\{b,d\},\{a,c,e\})$ by [3]. ∎
@@
We can now describe the facets of the matroids $\mathcal{M}_{T}$. We write $T\pm 1$ for the triangulation of the $m$-gon with triangles $H\pm 1$ where $H$ is a triangle of $T$.
@@
The facets of the oriented matroid $\mathcal{M}_{T}$ are given by $H$ where either $H$ is a triangle of $T$ or $H$ is a triangle of $T+1$.
We refer to facets $H$ where $H$ is a triangle of $T$ as upper facets of $\mathcal{M}_{T}$ and facets $H$ where $H$ is a triangle of $T+1$ as lower facets. This follows standard practice for cyclic polytopes; see [7]. In order to prove Proposition 3.11, we will need to use [2, Lemma 3.5.8], which we reproduce for the convenience of the reader.
Lemma 3.12 ([2, Lemma 3.5.8]).
@@
where $X=(x_{2},\dots,x_{r})$ is an ordered basis of the hyperplane $E\setminus\underline{D}$ of $\mathcal{M}$. Then $s(e,f)$ does not depend upon the choice of $X$.
@@
We first show that every hyperplane given by a triangle $H$ of $T$ is a facet. Let $H=abc$ be a triangle of $T$. We must show that every $x\in[m]\setminus H$ has the same sign in this cocircuit. If $c\prec x\prec a$, then $\chi_{T}(x,a,b,c)=+1$ due to the arc $ac$ of $T$. If $a\prec x\prec b$, then $\chi_{T}(x,a,b,c)=-\chi_{T}(a,x,b,c)=+1$ due to the arc $ab$ of $T$. One can similarly verify for $b\prec x\prec c$ that we have $\chi_{T}(x,a,b,c)=+1$. Using Lemma 3.12, we conclude that the signatures of $m]\setminus H$ are $([m]\setminus H,\varnothing)$ and $(\varnothing,[m]\setminus H)$, and so $H$ is a facet of $\mathcal{M}_{T}$.
Now let $H=abc$ be such that H1H-1 is a triangle of $T$. If $x\in[m]\setminus H$ has $c\prec x\prec a$, then $\chi_{T}(x,a,b,c)=-1$ due to the arc (a1)(c1)(a-1)(c-1) of $T$. Proceeding similarly to above, it is clear that $x$ must always have the sign 1-1 whenever it lies in $m]\setminus H$. We conclude that $H$ is a facet of $\mathcal{M}_{T}$ in this case too.
We now show that all facets of $\mathcal{M}_{T}$ are either upper facets or lower facets. Indeed, suppose that abcabc is a facet of $\mathcal{M}_{T}$. By rotating, we may assume that $c=m$ and $a\neq 1$. Suppose that $\chi(1,a,b,c)=+1$. Then there must be an arc $xc$ of $T$ with $a\preccurlyeq x\prec b$. If $x\neq a$, then $\chi(x,a,b,c)=-\chi(a,x,b,c)=-1$ due to the arc $xc$, which contradicts the fact that abcabc is a facet. We conclude that $ac$ is an arc of $T$. One can show in a similar way that $ab$ and $bc$ are arcs of $T$ if they are not boundary edges of the $m$-gon, so that abcabc is an upper facet. A similar argument shows that if $\chi(1,a,b,c)=-1$, then abcabc is a lower facet. Hence, all facets of $\mathcal{M}_{T}$ are either upper facets or lower facets. ∎
@@
Let $\sigma=\{a,b,c,d\}$ be a basis of $\mathcal{M}_{T}$ with $a\prec b\prec c\prec d$. Then if $\chi_{T}(a,b,c,d)=+1$, we say that $\{a,b,d\}$ and $\{b,c,d\}$ are upper facets of $\sigma$ with respect to $T$ and $\{a,b,c\}$ and $\{a,c,d\}$ are lower facets of $\sigma$ with respect to $T$.
@@
Suppose for contradiction that $\sigma\cap\tau$ is a lower facet of both $\sigma$ and $\tau$, with the case where it is an upper facet of both being similar. By Proposition 3.6, it suffices to show the claim when restricting to $\sigma\cup\tau$. However, $|\sigma\cup\tau|=5$, whence the restriction $\mathcal{M}_{T}(\sigma\cup\tau)$ is the oriented matroid of three-dimensional cyclic polytope with five vertices by Proposition 3.9 and [3]. In this case, the notions of upper and lower facets coincide with the usual notions for simplices in cyclic polytopes; see, for instance, [7, Section 2]. For cyclic polytopes, it follows for geometric reasons that if $\sigma\cap\tau$ is a lower facet of both $\sigma$ and $\tau$ then their geometric realisations $\sigma||$ and $\tau||$ intersect in more than this facet, since points a small distance above $\sigma||$ and $\tau||$ must lie in both $\sigma||$ and $\tau||$. It follows from this that there must be a circuit $C=(C^{+},C^{-})$ such that $\sigma\supseteq C^{+}$ and $\tau\supseteq C^{-}$. In fact, since $|\underline{C}|=5=|\sigma\cup\tau|$ and $|\sigma|=|\tau|=4$, if we let $\sigma\setminus\tau=\{a\}$, then it follows that $a\in C^{+}$ and $\underline{C}\setminus\{a\}\subseteq\tau$. Hence, $\sigma$ and $\tau$ overlap on a circuit, meaning that they cannot be elements of the same triangulation, and so we have arrived at a contradiction. ∎
@@
Given a triangulation $\Delta$ of $\mathcal{M}_{T}$, we define a relation $\rightsquigarrow$ on $\Delta$ to be the smallest reflexive transitive relation such that $\sigma\rightsquigarrow\tau$ whenever $\sigma\cap\tau$ is an upper facet of $\sigma$ and a lower facet of $\tau$. Following [19, Definition 2.13], we say that a triangulation $\Delta$ of $\mathcal{M}_{T}$ is stackable if the relation $\rightsquigarrow$ is a partial order on $\Delta$.

### v3

@@ 3. Constructing the oriented matroids
For now, let $T$ be a triangulation of a convex $m$-gon, where $m\geqslant 4$. We label the vertices of the $m$-gon anti-clockwise by the elements of $[m]$. We construct an oriented matroid $\mathcal{M}_{T}$ from $T$ by specifying its chirotope. Given $x,y\in[m]$ with $x\neq y$, we denote the arc with vertices $x$ and $y$ by $xy$. We refer to a subset of $[m]$ of size $k$ as a $k$-subset.
@@
If the triangulation $T$ contains arcs $xy$ with $a\preccurlyeq x\prec y\prec b$, then $T$ restricts to a triangulation of the polygon with vertices $V=[m]\setminus\{v\in[m]:x\prec v\prec y\}$, where we also have $V\supseteq\{a,b,c,d\}$. Hence, we can assume that there are no such arcs $xy$ of $T$, and likewise assume that $T$ contains no arcs $xy$ with $b\preccurlyeq x\prec y\prec c$, $c\preccurlyeq x\prec y\prec d$, or $d\preccurlyeq x\prec y\prec a$. Thus, if $x$ is a vertex with $a\preccurlyeq x\prec b$, then there must be an arc (or boundary edge) $xy$ of $T$ with either $b\preccurlyeq y\prec c$, $c\preccurlyeq y\prec d$, or $d\preccurlyeq y\prec a$. If the middle case holds, then we are done, so we can assume ($\ast$) that the first or the third case holds for all $x$ with $a\preccurlyeq x\prec b$. If we have $a\preccurlyeq x\prec x^{\prime}\prec b$, with $x^{\prime}y^{\prime}$ an arc of $T$, then we cannot have $b\preccurlyeq y\prec c$ and $d\preccurlyeq y^{\prime}\prec a$, since then $xy$ and $x^{\prime}y^{\prime}$ cross. Thus, let $\overrightarrow{x}$ be the maximal vertex with $a\preccurlyeq\overrightarrow{x}\prec b$ such that there is an arc $\overrightarrow{x}\overrightarrow{y}$ of $T$ with $d\preccurlyeq\overrightarrow{y}\prec a$. Similarly, let $\overleftarrow{x}$ be the minimal vertex with $a\preccurlyeq\overleftarrow{x}\prec b$ such that there is an arc $\overleftarrow{x}\overleftarrow{y}$ of $T$ with $b\preccurlyeq\overleftarrow{y}\prec c$. By assumption ($\ast$), there are no vertices $x$ with $\overrightarrow{x}\prec x\prec\overleftarrow{x}$. In fact, one can see that $\overrightarrow{x}=\overleftarrow{x}$, since one can otherwise consider the triangle with the side $\overrightarrow{x}\overleftarrow{x}$, which must either contain an arc $\overrightarrow{x}y$ with $b\preccurlyeq y\prec c$ or an arc $\overleftarrow{x}y$ with $d\preccurlyeq y\prec a$.
We thus let $x_{ab}=\overrightarrow{x}=\overleftarrow{x}$. One can deduce the existence of analogous vertices $x_{bc}$, $x_{cd}$, and $x_{da}$ with $b\preccurlyeq x_{bc}\prec c$, $c\preccurlyeq x_{cd}\prec d$, and $d\preccurlyeq x_{da}\prec a$. We have that there are arcs $\overleftarrow{y}x_{ab}$ and $x_{ab}\overrightarrow{y}$ with $d\preccurlyeq\overleftarrow{y}\prec a$ and $b\preccurlyeq\overrightarrow{y}\prec a$, and analogous arcs for each of $x_{bc}$, $x_{cd}$, and $x_{da}$. If none of these arcs are to intersect, we must have that they form a quadrilateral with vertices $x_{ab}$, $x_{bc}$, $x_{cd}$, and $x_{da}$, so that in particular $\overleftarrow{y}=x_{da}$ and $\overrightarrow{y}=x_{bc}$. The diagonal $xy$ of this quadrilateral which is an arc of $T$ then gives us the arc in the statement of the lemma. ∎
We now need to show that $\mathcal{M}_{T}$ is a well-defined oriented matroid. We will use [BLVS+99, Corollary 3.6.3], which says that $\mathcal{M}_{T}$ is well-defined if $\chi_{T}$ is alternating, the underlying (unoriented) matroid of $\mathcal{M}_{T}$ is well-defined, and every six-element restriction of $\mathcal{M}_{T}$ is realisable. The first property we have required to be true by construction. The second property is immediate, since the underlying matroid of $\mathcal{M}_{T}$ is the uniform matroid. To show the final property, we need to understand restrictions of $\mathcal{M}_{T}$.
@@
Given an arc $xy$ which separates $V$, set $x^{V}:=\max_{\circlearrowleft}V_{x}$ and $y^{V}:=\max_{\circlearrowleft}V_{y}$, where $\max_{\circlearrowleft}V_{x}$ is defined to be the unique element of $V_{x}$ such that $v\preccurlyeq\max_{\circlearrowleft}V_{x}\preccurlyeq x$ for all elements $v\in V_{x}$, and $\max_{\circlearrowleft}V_{y}$ is defined similarly. Define $T_{V}$ to be the triangulation of the polygon with vertex set $V$ whose set of arcs is
$\{x^{V}y^{V}:xy\text{ an arc of $T$ which separates }V\}$
[…]
@@
Note first that if $|V|=m$, then the statement is immediate, since in that case all arcs of $T$ separate $V$. It suffices then to consider the case where |V|=m1|V|=m-1, since one can use induction to obtain the statement for smaller $V$. We can then rotate to assume that $V=[m]\setminus\{m\}$. In this case, $T_{V}$ has arcs
@@
[…]
The fact that this is a triangulation of the convex polygon with vertices $V$ then follows from [Ram97, Theorem 4.2(iii)], but is also not hard to see by inspection. ∎
@@
Let $\{a\prec b\prec c\prec d\}\subseteq V$. Then $\chi_{T}(a,b,c,d)=+1$ if and only if there is an arc $xy$ of $T$ with $a\prec b\preccurlyeq x\prec c\prec d\preccurlyeq y$. Note that $xy$ therefore separates $V$. This is then the case if and only if $a\prec b\preccurlyeq x^{V}\prec c\prec d\preccurlyeq y^{V}$. Finally, this is the case if and only if there exists an arc $x^{V}y^{V}$ of $T_{V}$ such that this holds, which is the condition for $\chi_{T_{V}}(a,b,c,d)=+1$. ∎
Hence, using [BLVS+99, Corollary 3.6.3], in order to show that $\chi_{T}$ is a well-defined chirotope, it suffices to show that $\chi_{T}$ is realisable for $T$ a triangulation of a hexagon.
@@
If $T$ is the triangulation $\{26,36,46\}$, then one can check that $\chi_{T}$ is positive on all 4-subsets of $[6]$ ordered in the usual way. Hence, $\chi_{T}$ is realised by the three-dimensional cyclic polytope with six vertices from Example 2.3. By rotating, we have that $\chi_{T}$ is realisable for any triangulation $T$ of the form $\{(a-4)a,(a-3)a,(a-2)a\}$.
@@
Note that Table 1 gives examples of the the oriented matroids defined in Definition 3.1, via taking the signs of the minors. We then obtain the following result from Proposition 3.7 by applying [BLVS+99, Corollary 3.6.3].
@@
First, note that every 5-subset of $[m]$ is a circuit of $\mathcal{M}_{T}$, since every 4-subset is a basis. Thus, let $\underline{C}=\{a\prec b\prec c\prec d\prec e\}\subseteq[m]$ be the underlying set of a signed circuit of $\mathcal{M}_{T}$, where $a\prec b\prec c\prec d\prec e$. The signs of $C$ are determined by the restriction $\mathcal{M}_{T}(\underline{C})=\mathcal{M}_{T_{\underline{C}}}$ by Proposition 2.4. We have that $T_{\underline{C}}$ is a triangulation of a pentagon. Up to rotation, there is only one triangulation of a pentagon, so we can rotate and assume that $T_{\underline{C}}=\{be,ce\}$. We then have that $\mathcal{M}_{T_{\underline{C}}}$ is then the oriented matroid of a three-dimensional cyclic polytope with five vertices. This implies that $C=(\{a,c,e\},\{b,d\})$ or $C=(\{b,d\},\{a,c,e\})$, as in Example 2.3. ∎
@@
Since $\mathcal{M}_{T}$ is acyclic, we may talk about the facets of the matroids $\mathcal{M}_{T}$. We divide these into two types: upper facets and lower facets. This follows standard practice for cyclic polytopes; see [ER96].
Definition 3.11.
A facet abcabc is an upper facet of $\mathcal{M}_{T}$ if and only if $\chi_{T}(x,a,b,c)=+1$ for all $x\in[m]\setminus\{a,b,c\}$. Similarly, a facet abcabc is a lower facet of $\mathcal{M}_{T}$ if and only if $\chi_{T}(x,a,b,c)=-1$ for all $x\in[m]\setminus\{a,b,c\}$.
It follows from Lemma 3.13 that all facets of $\mathcal{M}_{T}$ are either upper facets or lower facets. We can then describe the upper and lower facets of $\mathcal{M}_{T}$ as follows. We write $T\pm 1$ for the triangulation of the $m$-gon with triangles $H\pm 1$ where $H$ is a triangle of $T$.
@@
The upper facets of the oriented matroid $\mathcal{M}_{T}$ are given by $H$ where $H$ is a triangle of $T$, and the lower facets are given by $H$ where $H$ is a triangle of $T+1$.
In order to prove Proposition 3.12, we will need to use [BLVS+99, Lemma 3.5.8], which we reproduce for the convenience of the reader. This lemma tells us how the chirotope determines the signatures of the cocircuits.
Lemma 3.13 ([BLVS+99, Lemma 3.5.8]).
@@
where $X=(x_{2},\dots,x_{r})$ is an ordered basis of the hyperplane $E\setminus\underline{D}$ of $\mathcal{M}$. (Here, a basis of the hyperplane is a maximally independent set contained in the hyperplane.) Then $s(e,f)$ does not depend upon the choice of $X$.
@@
In our case, the underlying unoriented matroid of $\chi_{T}$ is the rank-four uniform matroid on $[m]$, so the bases are precisely the 4-subsets of $[m]$. This means that the (unsigned) circuits are precisely the 5-subsets of $[m]$. The (unsigned) cocircuits are precisely the subsets of $[m]$ which do not intersect any circuit in exactly one element, and so are the (m3)(m-3)-subsets of $[m]$. As in Definition 2.5, the hyperplanes are the complements of the cocircuits, and so are exactly the $3$-subsets of $[m]$. This final fact will be particularly useful to bear in mind in the following proof. We will use Lemma 3.13 to show that the $3$-subsets described in Proposition 3.12 are the complements of positive cocircuits, which gives that they are facets by Definition 2.5.
@@
We first show that every hyperplane given by a triangle $H$ of $T$ is an upper facet. Let $H=abc$ be a triangle of $T$. We must show that every $x\in[m]\setminus H$ has the same sign in this cocircuit. If $c\prec x\prec a$, then $\chi_{T}(x,a,b,c)=+1$ due to the arc $ac$ of $T$. If $a\prec x\prec b$, then $\chi_{T}(x,a,b,c)=-\chi_{T}(a,x,b,c)=+1$ due to the arc $ab$ of $T$. One can similarly verify for $b\prec x\prec c$ that we have $\chi_{T}(x,a,b,c)=+1$. We conclude that abcabc is an upper facet of $\mathcal{M}_{T}$. Moreover, using Lemma 3.13, one can see that the signatures of $m]\setminus H$ are $([m]\setminus H,\varnothing)$ and $(\varnothing,[m]\setminus H)$, and so $H$ is indeed a facet of $\mathcal{M}_{T}$.
Now let $H=abc$ be such that H1H-1 is a triangle of $T$. If $x\in[m]\setminus H$ has $c\prec x\prec a$, then $\chi_{T}(x,a,b,c)=-1$ due to the arc (a1)(c1)(a-1)(c-1) of $T$. Proceeding similarly to above, it is clear that $x$ must always have the sign 1-1 whenever it lies in $m]\setminus H$. We conclude that $H$ is a lower facet of $\mathcal{M}_{T}$.
We now show that all facets of $\mathcal{M}_{T}$ are given by $H$ for $H$ either a triangle of $T$ or a triangle of $T+1$. Indeed, suppose that abcabc is an upper facet of $\mathcal{M}_{T}$. By rotating, we may assume that $c=m$ and $a\neq 1$. Since abcabc is an upper facet, we have $\chi(1,a,b,c)=+1$. Then there must be an arc $xc$ of $T$ with $a\preccurlyeq x\prec b$. If $x\neq a$, then $\chi(x,a,b,c)=-\chi(a,x,b,c)=-1$ due to the arc $xc$, which contradicts the fact that abcabc is a facet. We conclude that $ac$ is an arc of $T$. One can show in a similar way that $ab$ and $bc$ are arcs of $T$ if they are not boundary edges of the $m$-gon, so that abcabc is an upper facet. A similar argument shows that if abcabc is a lower facet, then it is a triangle of $T+1$. Hence, all facets of $\mathcal{M}_{T}$ are given by triangles of $T$ or $T+1$. ∎
@@
Let $\sigma=\{a,b,c,d\}$ be a basis of $\mathcal{M}_{T}$ with $a\prec b\prec c\prec d$. Then if $\chi_{T}(a,b,c,d)=+1$, we say that $\{a,b,d\}$ and $\{b,c,d\}$ are upper facets of $\sigma$ with respect to $T$ and $\{a,b,c\}$ and $\{a,c,d\}$ are lower facets of $\sigma$ with respect to $T$. If $\chi_{T}(a,b,c,d)=-1$, then $\chi_{T}(d,a,b,c)=+1$, so the definition reduces to the former case.
@@
Suppose for contradiction that $\sigma\cap\tau$ is a lower facet of both $\sigma$ and $\tau$, with the case where it is an upper facet of both being similar. By Proposition 3.6, it suffices to show the claim when restricting to $\sigma\cup\tau$. However, $|\sigma\cup\tau|=5$, whence the restriction $\mathcal{M}_{T}(\sigma\cup\tau)$ is the oriented matroid of a three-dimensional cyclic polytope with five vertices by Proposition 3.9. In this case, the notions of upper and lower facets coincide with the usual notions for simplices in cyclic polytopes; see, for instance, [ER96, Section 2]. For cyclic polytopes, if $\sigma\cap\tau$ is a lower facet of both $\sigma$ and $\tau$ then their geometric realisations $\sigma||$ and $\tau||$ intersect in their interiors, since points a small distance above $\sigma||$ and $\tau||$ must lie in both $\sigma||$ and $\tau||$. (Here we are using the standard realisation of cyclic polytopes as the convex hull of a finite set of points on the moment curve — see for instance [Wil22, Section 2.1] — and $\sigma||$ denotes the geometric realisation of $\sigma$ as the convex hull of its vertex set on the moment curve.)
Since $\sigma||$ and $\tau||$ intersect in their interiors, there is therefore a circuit $C=(C^{+},C^{-})$ such that $\sigma\supseteq C^{+}$ and $\tau\supseteq C^{-}$. (In other words, either $\sigma$ has a one-dimensional face which intersects a two-dimensional face of $\tau$ or vice versa.) In fact, since $|\underline{C}|=5=|\sigma\cup\tau|$ and $|\sigma|=|\tau|=4$, we have $|\sigma\setminus\tau|=1$. We can then let $\sigma\setminus\tau=\{a\}$, and so $a\in C^{+}$ and $\underline{C}\setminus\{a\}\subseteq\tau$. Hence, $\sigma$ and $\tau$ overlap on a circuit, meaning that they cannot be elements of the same triangulation, and so we have arrived at a contradiction. ∎
@@
Given a triangulation $\Delta$ of $\mathcal{M}_{T}$, we define a relation $\rightsquigarrow$ on $\Delta$ to be the smallest reflexive transitive relation such that $\sigma\rightsquigarrow\tau$ whenever $\sigma\cap\tau$ is an upper facet of $\sigma$ and a lower facet of $\tau$. Following [RS00, Definition 2.13], we say that a triangulation $\Delta$ of $\mathcal{M}_{T}$ is stackable if the relation $\rightsquigarrow$ is a partial order on $\Delta$. It suffices to show that $\rightsquigarrow$ is an anti-symmetric relation.